Circular No. 8088

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

S/2003 J 8 Further to IAUC 8087, S. S. Sheppard reports another satellite of

Jupiter found on CCD images taken at Mauna Kea during Feb. 8–Mar. 5 (MPEC 2003-E15). After computing the preliminary orbit, B. G. Marsden made a linkage to Mauna Kea observations in Jan. 2002, then finding e = 0.26, i = 153^{o}, P = 781 days, H = 15.9 (MPEC 2003-E24).

SUPERNOVA 2003bn Further to IAUC 8084, W. M. Wood-Vasey, G. Aldering, and P. Nu-

gent report the discovery of an apparent supernova (mag 16.7) on unfiltered NEAT images taken with the Haleakala 1.2-m telescope on 2003 Feb. 27.48 and 28.38 UT. The new object is located at α = 10^{h}02^{m}35^{s}.51, δ = −21^{o}10′54′′.5 (equinox 2000.0), which is 3′′.2 east and 1′′.0 south of the center of the apparent host galaxy. Additional magnitudes from NEAT images: Feb. 21.0, [21.0; 22.33, 20.2. J. Mulchaey and M. Phillips, Las Campanas Observatory, found SN 2003bn at V ∼ 16.7 on Mar. 6.13.

SUPERNOVA 2003bg IN MCG –05-10-15 M. Hamuy, M. Phillips, and J. Thomas-Osip, Carnegie Observatories,

write: “An optical spectrum of SN 2003bg, obtained on Mar. 4 with the Las Campanas Observatory 2.5-m telescope, indicates the presence of strong, broad Hα emission with a P-Cyg absorption component with an expansion velocity of −17110 km/s at its minimum. P-Cyg absorption minima apparently corresponding to Hβ and Hγ are also present at expansion velocities of −14600 and −13650 km/s. Although the overall appearance of the spectrum of SN 2003bg is similar to that of other broad-lined type-Ic events, the presence of strong hydrogen features distinguishes it.”

IGR J19140+098 D. C. Hannikainen, University of Helsinki Observatory; J. Rodriguez,

CEA, Service d’Astrophysique, Saclay, and Integral Science Data Center (ISDC), Versoix; and K. Pottschmidt, Max Planck Institut für Extraterrestrische Physik, Garching, and ISDC, report the detection by INTE-GRAL (+ IBIS/ISGRI) on Mar. 6.456 UT (during an observation of GRS 1915+105), of a new transient at α = 19^{h}13^{m}55^{s}, δ = +9^{o}51′.6 (equinox 2000.0; uncertainty 1′). The source, which varied significantly during a span of 2000 s, had a flux in the band 15–40 keV of 50–100 mCrab.

2003 March 6  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2003 CBAT  Daniel W. E. Green